Multi-Variable Calculus Assignment

due Friday, February 23, 2001

Exercises

1. Find the equation of the plane tangent at the point (1, 1, 1) to the surface xyz − x^{2}y + y^{2}z − z^{2}x = 0 .

2. Find the equation of the plane normal at the point (1, −1, 1) to the curve in space given

parametrically by

t ⟼ (t − 1, t^{4} − 3t^{3} + t^{2} + 3, t^{3} − 3t − 1) .

[Processed from GELLMU to L^{A}TEX on Wed 21 Feb 2001 at 08:58:07 EST]